OpenStudy (anonymous):

Substitution - add/subtraction method for 2 equations: x + 3y = 19 x - y = -1

OpenStudy (anonymous):

First, just isolate one of the variables, Let's isolate (x): x - y = -1 -y = -1 - x y = 1 + x now lets plug in the (y) into the other equation: x + 3y = 19 x + 3(1 + x) = 19 x + 3 + 3x = 19 4x + 3 = 19 4x = 16 x = 4 now lets use this to find our (y) y = 1 + x y = 1 + 4 y = 5 so x = 4, y = 5